OpenStudy (maddiie):

In a paragraph, describe how to simplify each expression. 16 1/2 Square root 100

OpenStudy (anonymous):

????????? why don't u use the equation box to write yr question...

OpenStudy (maddiie):

In a paragraph, describe how to simplify each expression. 16 1/2 \[\sqrt{100}\]

OpenStudy (anonymous):

is the first one \[16^{1/2}\] ???

OpenStudy (maddiie):

yes

OpenStudy (anonymous):

pls note that \[x ^{1/2} = \sqrt{x}\]

OpenStudy (maddiie):

why is that

OpenStudy (anonymous):

okay I will explain thru an example...

OpenStudy (maddiie):

ok awesome

OpenStudy (anonymous):

u know that \[\sqrt{4}= 2\] also \[(4)^{1/2} = (2^{2})^{1/2} = 2^{2*1/2} = 2^{1}=2\]

OpenStudy (anonymous):

so we find that \[\sqrt{4} = (4)^{1/2}\] because both give same result i.e. 2

OpenStudy (maddiie):

\[({8}^2)\]

OpenStudy (maddiie):

confused

OpenStudy (anonymous):

so we can solve the first one as \[(16)^{1/2} = (4^{2})^{1/2} = 4^{2*1/2} = 4^{1}=4\]

OpenStudy (anonymous):

what is confusing u???

OpenStudy (maddiie):

Oh because 4x4=16 right

OpenStudy (anonymous):

yes

OpenStudy (maddiie):

So \[\sqrt{100}= 10 \]

OpenStudy (anonymous):

You might want to introduce and include rules of exponents and roots as part of your explanation.

OpenStudy (anonymous):

yes that is correct

OpenStudy (maddiie):

ok i understood the steps now

OpenStudy (maddiie):

ok i understood the steps now

OpenStudy (anonymous):

see square and root are opposite operations. so when they occur together, they cancel each other out.... we can show the working for root100 as follows : \[\sqrt{100}=\sqrt{10^{2}}=10\] the square inside and root outside cancel each other out

OpenStudy (anonymous):

are u familiar with exponential laws???

OpenStudy (maddiie):

makes sense

OpenStudy (maddiie):

not really but i can try to find it in my book

OpenStudy (anonymous):

yes, u shud have all the laws in front of you and slowly try to learn them by heart....

OpenStudy (maddiie):

ok thank u

OpenStudy (anonymous):

u r welcome, just keep in mind... \[\sqrt{x}=x ^{1/2}\] \[\sqrt[3]{x}=x ^{1/3}\] \[\sqrt[4]{x}=x ^{1/4}\] and so on....